Suppose LaTeX: S=\frac{d}{1+k}+\frac{d(1+g)}{(1+k)^2}+\frac{d(1+g)^2}{(1+k)^3}+...+\frac{d(1+g)^n}{(1+k)^{n+1}} , obtain the value of S if d=7, g=0.04, k=0.07, and n=22. Round your final answer to two decimal places.
Answer #1

S = D/(1+k) + d*(1+g)/(1+k)2 + d*(1+g)2/(1+k)3 + ......d*(1+g)n/(1+k)n+1

= 7/(1+0.07) + 7*(1+0.04)/(1+0.07)2 + 7*(1+0.04)2/(1+0.07)​​​​​​3.......+ 7*(1.04)22/1.0723 = 112.02

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well